# Counting spaces of functions on separable compact lines

### v1

@@ 5. General properties of spaces $C(K_{A})$
The next result follows from a classical theorem of Heath and Lutzer [12, Theorem 2.4] which provides the existence of an extension operator $T:C(L)\to C(K)$ for compact lines $L\subseteq K$ (see [21, Lemma 4.3] for a self-contained proof).
@@
The following result is a more general version of [21, Lemma 4.6]. Note that it is true in ZFC, without any additional set-theoretical assumptions. The proof of this lemma follows closely the argument in [21], we present it here for the convenience of readers.
@@
i.e., a positive linear operator $T:C(C_{B})\to C(C_{1})$ satisfying $T(1_{C_{B}})=1_{C_{1}}$ and $T(g\circ\varphi)=g$ for every $g\in C(C_{1})$. The existence of such an operator implies that $C(C_{1})|C(L)$, see [30, Proposition 8.2].
@@
for $f\in C(C_{B})$ and $x\in C_{1}$. It is a well-known fact (see e.g. [21, Proof of Thm. 3.7]) that, for every function $f\in C(C_{B})$, the function $f_{0}(t)=f(t,0)$ has at most countably many points of discontinuity and is therefore measurable. It follows that $T(f)(x)$ is well defined for every $x\in C_{1}$, it is also clear that $T$ is positive and $T(1_{C_{B}})=1_{C_{1}}$. Additionally, for every $g\in C(C_{1})$, $x\in C_{1}$ and $y\in C_{2}$ we have $g\circ\varphi(h^{-1}(x,y),0)=g(x)$, hence $T(g\circ\varphi)=g$.
It remains to check that $T(f)$ is continuous on $C_{1}$ for every $f\in C(C_{B})$. Consider any converging sequence $x_{n}\to x$ in $C_{1}$. Then for every $y\in C_{2}$ we have $h^{-1}(x_{n},y)\to h^{-1}(x,y)$, as $h$ is a homeomorphism. Consider again the function $f_{0}(t)=f(t,0)$ and let $P$ be the set of points of discontinuity of $f_{0}$ ($P$ is countable). Denote $f_{n}(y)=f(h^{-1}(x_{n},y),0)$. If $R=\pi_{2}\circ h(P)$, then for every point $y\in C_{2}\setminus R$ we have $f_{n}(y)\to f(y)$, thus the sequence $(f_{n})_{n\in\omega}$ is $\mu$-a.e. convergent to $f$ on $C_{2}$ (note that $\mu(R)=0$, as $R$ is countable and the measure $\mu$ is nonatomic). As all functions $|f_{n}|$ are bounded by the norm of $f$, from the Lebesgue dominated convergence theorem it follows that $T(f)(x_{n})\to T(f)(x)$.
@@
$C(L)\simeq C(L^{(\alpha)})\oplus C_{0}(L\|L^{(\alpha)})\simeq C(L^{(\alpha)})\oplus C(K)\simeq C(L^{(\alpha)}).$
@@
Let $L$ be an uncountable separable compact line. Then there exists $A\subset(0,1)$ such that $C(L)\simeq C(I_{A})$.
@@ 6. Counting spaces $C(K_{A})$ of weight $\kappa$ when $2^{\kappa}>2^{\omega}$
It is a classical fact that for any separable compact line $K_{A}$ we have $|M(K_{A})|=2^{\omega}$, see [17].
@@ 7. Counting spaces $C(K_{A})$ under Baumgartner’s Axiom
We will mainly use this definition in the case when $X$ is a subspace of the real line $\mathbb{R}$, so we will consider only cardinals $\kappa\leq 2^{\omega}$. Obviously, a subset $A$ of $\mathbb{R}$ is $\kappa$-dense in $\mathbb{R}$ if, for every nonempty open interval $I\subseteq\mathbb{R}$ we have $|I\cap A|=\kappa$. We refer to the following statement for $\kappa$-dense sets in $\mathbb{R}$ as Baumgartner’s Axiom
BA($\kappa$): Every two sets of reals, which are $\kappa$-dense in $\mathbb{R}$, are order isomorphic.
@@
Repeating the construction of the reals using the Dedekind cuts, one can easily obtain the following lemma.
@@
If, for $\omega_{1}$-dense sets $A,B\subseteq(0,1)$, there is an order isomorphism $f:A\to B$, then the spaces $I_{A},I_{B}$ are homeomorphic.
@@
The space $C(I_{A\cap U})$ is separable, since $A\cap U$ is countable; therefore its subspace $C_{0}(I_{A\cap U}\|M)$ is also separable (here, we consider $M$ as a subspace of $I_{A\cap U}$). Let $\varphi:I_{A}\to I_{A\cap U}$ be a natural continuous surjection defined by the formula:
@@
One can easily check that the isometric embedding $T_{\varphi}:C(I_{A\cap U})\to C(I_{A})$ generated by $\varphi$, i.e., an operator sending each $f\in C(I_{A\cap U})$ to the composition $f\circ\varphi$, maps the space $C_{0}(I_{A\cap U}\|M)$ onto $C_{0}(I_{A}\|M_{A\setminus U})$, hence the last space is also separable.
By Corollary 5.2 the space $C_{0}(K\|L)$ is isomorphic to the space $C(N)$ for some compact space $N$. Since $C_{0}(K\|L)$ is separable, $N$ must be metrizable, so we can use Lemma 5.4 and property (1) to conclude that $C(K)$ and $C(L)$ are isomorphic.
@@
$C(K)\simeq C(L)\simeq C(L^{(\alpha)})\simeq C(I_{B})\,.$
@@ 8. Function spaces on finite products of separable compact lines
Using Theorem 1.3, the results of Michalak from [23], and assuming the Baumgartner’s Axiom we obtain, consistently, a complete isomorphic classification of spaces $C(K)$, where $K$ is a finite product of separable compact lines of weight $\leq\omega_{1}$.
@@
$\cup\{(I_{\mathbf{A}})^{n}\times(\omega^{\omega^{\alpha}}+1):n\geq 1,0<\alpha<\omega_{1}\}.$
The goal of this Section is to prove that the class $\mathcal{K}_{\mathbf{A}}$ (regardless of which set $A\subseteq(0,1)$ of cardinality $\omega_{1}$ is chosen) exhausts all the isomorphism types of $C(K)$, where $K$ is any finite product of separable compact lines of weight $\leq\omega_{1}$. First, let us note that the spaces $C(K)$ for $K\in\mathcal{K}_{\mathbf{A}}$ are pairwise nonisomorphic due to the following result of Michalak.
@@
First, recall a simple observation that, for any compact space $K$, the space $C(K,c_{0})$ is isometric to the $c_{0}$-sum of $\omega$ copies of $C(K)$. Indeed, $c_{0}$ is isometric to $C_{0}(\omega)$, hence $C(K,c_{0})$ is isometric to $C(K,C_{0}(\omega))$, which in turn is isometric to $C_{0}(K\times\omega)$, see [32]*Proposition 7.7.6. Clearly, this last space is isometric to the $c_{0}$-sum of $\omega$ copies of $C(K)$.
@@
As mentioned previously, by Theorem 8.1, for two distinct spaces $K_{1},K_{2}\in\mathcal{K}_{\mathbf{A}}$, we have $C(K_{1})\not\simeq C(K_{2})$. Thus, it is enough to prove that, for any separable compact lines $K_{1},\dotsc,K_{m}$ of weight $\leq\omega_{1}$, if at least one of them is nonmetrizable, then there is some $L\in\mathcal{K}_{\mathbf{A}}$ such that $C(\prod_{i=1}^{m}K_{i})\simeq C(L)$.
Without loss of generality, assume that for some $n\leq m,n\geq 1$, we have $w(K_{i})=\omega_{1}\iff i\leq n$. Then $M=\prod_{i=n+1}^{m}K_{i}$ is a metrizable compact space.
@@
$C(\prod_{i=1}^{m}K_{i})\simeq C((I_{\mathbf{A}})^{n}\times M).$
@@
$C((I_{\mathbf{A}})^{n}\times M)\simeq C((I_{\mathbf{A}})^{n},C(M))\simeq C((I_{\mathbf{A}})^{n},C(M^{\prime}))\simeq C((I_{\mathbf{A}})^{n}\times M^{\prime}).$
@@
due to the fact that the space $I_{\mathbf{A}}\times M^{\prime}$ is homeomorphic to a compact line.
@@ 9.1. Compact spaces of weight $\kappa$
Theorem 1.1 should be compared with the corollary from the Miljutin, Bessaga and Pełczyński’s classification of separable Banach spaces of the form $C(K)$ which says that there are only $\omega_{1}$ mutually nonisomorphic such spaces regardless of the relation between $\omega_{1}$ and $2^{\omega}$ (although the spaces $\ell_{p}$ for $1<p<\infty$ form a family of mutually nonisomorphic separable Banach spaces of cardinality $2^{\omega}$). Our method of proving Proposition 1.1 works only for regular cardinals. In particular, we do not know the answer to the following:
@@
Note that the answer to the above questions is positive if we assume the Generalized Continuum Hypothesis (GCH). Indeed, Kisljakov’s classification of spaces of the form $C([0,\alpha])$ for $\alpha<\kappa^{+}$ ([18]) shows that without any additional hypothesis we have $\kappa^{+}$ mutually nonisomorphic Banach spaces of the form $C(K)$ for any uncountable cardinal $\kappa$. As GCH means that $2^{\kappa}=\kappa^{+}$ for every infinite cardinal $\kappa$ we obtain the positive answer to the above questions under GCH.

### v3

@@ 5. General properties of spaces $C(K_{A})$
The next result follows from a theorem of Heath and Lutzer [13, Theorem 2.4] which provides the existence of an extension operator $T:C(L)\to C(K)$ for compact lines $L\subseteq K$ (see [23, Lemma 4.3] for a self-contained proof).
@@
The following result is a more general version of [23, Lemma 4.6]. Note that it is true in ZFC, without any additional set-theoretic hypothesis. The proof of this lemma follows closely the argument in [23], we present it here for the convenience of readers.
@@
i.e., a positive linear operator $T:C(C_{B})\to C(C_{1})$ satisfying $T(1_{C_{B}})=1_{C_{1}}$ and $T(g\circ\varphi)=g$ for every $g\in C(C_{1})$. The existence of such an operator implies that $C(C_{1})|C(L)$, see [32, Proposition 8.2].
@@
for $f\in C(C_{B})$ and $x\in C_{1}$. It is a well-known fact that, for every function $f\in C(C_{B})$, the function $f_{0}(t)=f(t,0)$ has at most countably many points of discontinuity (see e.g. [23, Proof of Thm. 3.7]) and is therefore measurable. It follows that $T(f)(x)$ is well defined for every $x\in C_{1}$, it is also clear that $T$ is positive and $T(1_{C_{B}})=1_{C_{1}}$. Additionally, for every $g\in C(C_{1})$, $x\in C_{1}$ and $y\in C_{2}$ we have $g\circ\varphi(h^{-1}(x,y),0)=g(x)$, hence $T(g\circ\varphi)=g$.
It remains to check that $T(f)$ is continuous on $C_{1}$ for every $f\in C(C_{B})$. Consider any convergent sequence $x_{n}\to x$ in $C_{1}$. Then for every $y\in C_{2}$ we have $h^{-1}(x_{n},y)\to h^{-1}(x,y)$, as $h$ is a homeomorphism. Consider again the function $f_{0}(t)=f(t,0)$ and let $P$ be the set of points of discontinuity of $f_{0}$ ($P$ is countable). Denote $f_{n}(y)=f_{0}(h^{-1}(x_{n},y))$ and $\widetilde{f}(y)=f_{0}(h^{-1}(x,y))$. If $R=\pi_{2}\circ h(P)$, then for every point $y\in C_{2}\setminus R$ we have $f_{n}(y)\to\widetilde{f}(y)$, thus the sequence $(f_{n})_{n\in\omega}$ is $\mu$-a.e. convergent to $\widetilde{f}$ on $C_{2}$ (note that $\mu(R)=0$, as $R$ is countable and the measure $\mu$ is nonatomic). As all functions $|f_{n}|$ are bounded by the norm of $f$, from the Lebesgue dominated convergence theorem it follows that $T(f)(x_{n})\to T(f)(x)$.
@@
$C(L)\simeq C(L^{(\alpha)})\oplus C_{0}(L\|L^{(\alpha)})\simeq C(L^{(\alpha)})\oplus C(K)\simeq C(L^{(\alpha)}).\qed$
@@
Let $L$ be an uncountable separable compact line. Then there exists $A\subseteq(0,1)$ such that $C(L)\simeq C(I_{A})$.
@@ 6. Counting spaces $C(K_{A})$ of weight $\kappa$ when $2^{\kappa}>2^{\omega}$
It is a known fact that for any separable compact line $K_{A}$ we have $|M(K_{A})|=2^{\omega}$, see [18].
@@ 7. Counting spaces $C(K_{A})$ under Baumgartner’s Axiom
We will mainly use this definition in the case when $X$ is a subspace of the real line $\mathbb{R}$, so we will consider only cardinals $\kappa\leq 2^{\omega}$. Obviously, a subset $A$ of $\mathbb{R}$ is $\kappa$-dense in $\mathbb{R}$ if, for every nonempty open interval $I\subseteq\mathbb{R}$, we have $|I\cap A|=\kappa$. We refer to the following statement about $\kappa$-dense sets in $\mathbb{R}$ as Baumgartner’s Axiom.
BA($\kappa$): Every two sets of reals that are $\kappa$-dense in $\mathbb{R}$ are order isomorphic.
@@
Using Dedekind cuts, one can easily obtain the following lemma.
@@
If $\omega_{1}$-dense sets $A,B\subseteq(0,1)$ are order isomorphic, then the spaces $I_{A},I_{B}$ are homeomorphic.
@@
The space $C(I_{A\cap U})$ is separable, since $A\cap U$ is countable; therefore its subspace $C_{0}(I_{A\cap U}\|M\times\{0\})$ is also separable. Let $\varphi:I_{A}\to I_{A\cap U}$ be a natural continuous surjection defined by the formula:
@@
One can easily check that the isometric embedding $T_{\varphi}:C(I_{A\cap U})\to C(I_{A})$ generated by $\varphi$, i.e., an operator sending each $f\in C(I_{A\cap U})$ to the composition $f\circ\varphi$, maps the space $C_{0}(I_{A\cap U}\|M\times\{0\})$ onto $C_{0}(I_{A}\|M_{A\setminus U})$, hence the latter space is also separable.
By Corollary 5.2 the space $C_{0}(K\|L)=C_{0}(I_{A}\|M_{A\setminus U})$ is isomorphic to the space $C(N)$ for some compact space $N$. Since $C_{0}(K\|L)$ is separable, $N$ must be metrizable, so we can use Lemma 5.4 and property (1) to conclude that $C(K)$ and $C(L)$ are isomorphic.
@@
$C(K)\simeq C(L)\simeq C(L^{(\alpha)})\simeq C(I_{B})\,.\qed$
@@ 8. Function spaces on finite products of separable compact lines
In this section, using Theorem 1.3, the results of Michalak from [25], and assuming Baumgartner’s Axiom, we consistently obtain a complete isomorphic classification of spaces $C(K)$, where $K$ is a nonmetrizable finite product of separable compact lines of weight $\leq\omega_{1}$.
@@
$\cup\{(I_{\mathbf{A}})^{n}\times(\omega^{\omega^{\alpha}}+1):n\geq 1,\omega_{1}>\alpha>0\}.$
The goal of this Section is to prove that the class $\mathcal{K}_{\mathbf{A}}$ (regardless of which set $\mathbf{A}\subseteq(0,1)$ of cardinality $\omega_{1}$ is chosen) exhausts all isomorphism types of $C(K)$, where $K$ is any nonmetrizable finite product of separable compact lines of weight $\leq\omega_{1}$. First, let us note that the spaces $C(K)$ for $K\in\mathcal{K}_{\mathbf{A}}$ are pairwise nonisomorphic due to the following result of Michalak.
@@
First, recall a simple observation that, for any compact space $K$, the space $C(K,c_{0})$ is isometric to the $c_{0}$-sum of $\omega$ copies of $C(K)$. Indeed, $c_{0}$ is isometric to $C_{0}(\omega)$, hence $C(K,c_{0})$ is isometric to $C(K,C_{0}(\omega))$, which in turn is isometric to $C_{0}(K\times\omega)$, see [34]*Proposition 7.7.6. Clearly, this latter space is isometric to the $c_{0}$-sum of $\omega$ copies of $C(K)$.
@@
By Theorem 8.1, for two distinct spaces $K_{1},K_{2}\in\mathcal{K}_{\mathbf{A}}$, we have $C(K_{1})\not\simeq C(K_{2})$. Thus, it is enough to prove that, for any $m\geq 1$, and any separable compact lines $K_{1},\dotsc,K_{m}$ of weight $\leq\omega_{1}$, if at least one of them is nonmetrizable, then there is some $L\in\mathcal{K}_{\mathbf{A}}$ such that $C(\prod_{i=1}^{m}K_{i})\simeq C(L)$.
Without loss of generality, assume that for some $1\leq n\leq m$, we have $w(K_{i})=\omega_{1}\iff i\leq n$. Then $M=\prod_{i=n+1}^{m}K_{i}$ is a metrizable compact space.
@@
$C\left(\prod_{i=1}^{m}K_{i}\right)\simeq C\big((I_{\mathbf{A}})^{n}\times M\big).$
@@
$C\big((I_{\mathbf{A}})^{n}\times M\big)\simeq C\big((I_{\mathbf{A}})^{n},C(M)\big)\simeq C\big((I_{\mathbf{A}})^{n},C(M^{\prime})\big)\simeq C\big((I_{\mathbf{A}})^{n}\times M^{\prime}\big).$
@@
$C\big((I_{\mathbf{A}})^{n}\times M^{\prime}\big)$ $\simeq$ $C\big((I_{\mathbf{A}})^{n-1}\times I_{\mathbf{A}}\times M^{\prime}\big)\simeq C\big((I_{\mathbf{A}})^{n-1},C(I_{\mathbf{A}}\times M^{\prime})\big)$
$\simeq$ $C\big((I_{\mathbf{A}})^{n-1},C(I_{\mathbf{A}})\big)\simeq C((I_{\mathbf{A}})^{n})$
due to Theorem 1.3 and the fact that the space $I_{\mathbf{A}}\times M^{\prime}$ is homeomorphic to a separable compact line.
@@ 9.1. Compact spaces of weight $\kappa$
Theorem 1.1 should be compared with the corollary from the Miljutin, Bessaga and Pełczyński’s classification of separable Banach spaces of the form $C(K)$ which says that there are only $\omega_{1}$ mutually nonisomorphic such spaces regardless of the relation between $\omega_{1}$ and $2^{\omega}$ (although the spaces $\ell_{p}$ for $1<p<\infty$ form a family of cardinality $2^{\omega}$ consisting of mutually nonisomorphic separable Banach spaces). Our method of proving Proposition 1.1 works only for regular cardinals. In particular, we do not know the answer to the following:
@@
Note that the answer to the above questions is positive if we assume the Generalized Continuum Hypothesis (GCH). Indeed, Kisljakov’s classification of spaces of the form $C([0,\alpha])$ for $\alpha<\kappa^{+}$ ([19]) shows that without any additional hypothesis we have $\kappa^{+}$ mutually nonisomorphic Banach spaces of the form $C(K)$ for any uncountable cardinal $\kappa$. As GCH means that $2^{\kappa}=\kappa^{+}$ for every infinite cardinal $\kappa$, we obtain the positive answer to the above questions under GCH.